OpenStudy (anonymous):

PLZ HALP Which is an example of the commutative property? A. –13 + 27 = 27 + (–13) B. –21 + (11 + 13) = (–21 + 11) + 13 C. 4(–3 + 5) = 4(–3) + 4 ∙ 5 D. –(–8 + 3) = 8 + (–3)

OpenStudy (anonymous):

SOME ONE?

OpenStudy (anonymous):

PLZ IMA NEWB

OpenStudy (lifeisadangerousgame):

The commutative property states that a * b = b * a, which answer matches that description?

OpenStudy (anonymous):

A?

OpenStudy (lifeisadangerousgame):

Yep! All that happens is that the places of 27 and -13 are switched.

OpenStudy (anonymous):

hue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue huehue hue hue hue hue hue hue hue hue

OpenStudy (jango_in_dtown):

a+b=b+a a*b=b*a

OpenStudy (anonymous):

ok how do i give you a gold medel

OpenStudy (anonymous):

cause u so awsome

OpenStudy (anonymous):

also thx can you help me with another one?

OpenStudy (anonymous):

@LifeIsADangerousGame

OpenStudy (lifeisadangerousgame):

Click on "Best Response" on one of my comments. Close this question and open up another one and I'll try to help you out!

OpenStudy (anonymous):

ok i already did

OpenStudy (lifeisadangerousgame):

Tag me in it please.